A gas contains 16.5 g of nitrogen and 3.5 g of hydrogen.what is it's simplest formula
1
Expert's answer
2020-03-31T12:32:57-0400

n(N) = 16.5 / 14 = 1.179 mol

n(H) = 3.5 / 1 = 3.5 mol

3.5 : 1.179 = 3 : 1

The formula of a gas is NH3